OpenStudy (anonymous):

The quantity of a charge Q that passes through a surface at time t (seconds), is given by Q=t^3 -2t^2 +6t +2. The current is defined as the rate of change of the charges with respect to t. Find the current when t=2

OpenStudy (anonymous):

\[i(t)=\frac{ dq }{dt }=3t^2-4t+6\] \[i(2)=3(2)^2-4(2)+6=10A\]

OpenStudy (anonymous):

Apparently the answer is \[\frac{ 3 }{ 10\Pi } \] cm/s

OpenStudy (anonymous):

what's this unit?

OpenStudy (anonymous):

related rates